# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 4.3. Mechanism Design
$\Delta_{2}(f)=$ $\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\mathbf{B}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}\|_{2}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\|\mathbf{B}\mathbf{d}\|_{2}=\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$
4.4. Ads-BPC Algorithm
We aim to minimize the total variance of noise added, for the mechanism defined in Eq. 4. Because ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$, $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, the total variance of the noise is $\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$. Recall that we calculate the matrix $\mathbf{L}$ as $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$, where $\mathbf{B}$ is a diagonal matrix. Therefore, the total variance can be represented as,
$\operatorname{trace}(\mathbf{L}\mathbf{L}^{T})$ $=\operatorname{trace}(\mathbf{W}\mathbf{B}^{-1}(\mathbf{W}\mathbf{B}^{-1})^{T})$
$=\operatorname{trace}(\mathbf{W}^{T}\mathbf{W}(\mathbf{B}^{T}\mathbf{B})^{-1})$
$=n\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}.$
The last equation holds because $(\mathbf{W}^{T}\mathbf{W})[i,i]=n-i+1,1\leq i\leq n$.
From Theorem 2.5, to satisfy $\rho$-zCDP, it’s required that the sensitivity is bounded $\Delta_{2}(f)\leq\sqrt{2\rho}$. Because we want to achieve $\rho$-zCDP while minimizing the total variance, the problem becomes,
(5) $\min_{\sigma_{1}^{2},\cdots,\sigma_{n}^{2}}\quad$ ${n}\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}$
$s.t.\quad$ $\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\frac{r_{2}^{2}}{\sigma_{2}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}\leq 2\rho.$
The following theorem shows that matrix $\mathbf{B}$ derived above satisfies $\rho$-zCDP. The proof is shown in Appendix 8.2.
@@ 4.4. Ads-BPC Algorithm
Let $\mathbf{W}\in\mathbb{R}^{n\times n}$ be the prefix-sum matrix 4.1. Let $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the solution to problem 5, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, then ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ = $\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$ satisfies $\rho$-zCDP.
Problem (5) needs to be solved efficiently for the online service. In Appendix 8.3, we obtain the following results:

### v2

@@ 4.2 Mechanism Privacy Budget
$\Delta_{2}(f)=$ $\max_{({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}^{\prime})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}-\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}^{\prime}\|_{2}=\max_{\mathbf{d}\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{User}}}}}\|\mathbf{B}\mathbf{d}\|_{2}$
$=\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$ (3)
We define the privacy budget of the mechanism as
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{budget}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}):=\frac{\Delta^{2}_{2}(f)}{2}=\rho.$ (4)
4.3 Initialize the Scales
With the expressions for utility and budget of our mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ established in previous sections, we can now optimize the parameters $\sigma_{i}$ for various queries and objective functions. Before running the ads measurement system, advertisers need to provide: 1) The specific queries of interest, such as the number of conversions each day or the number of conversions from day 10 to day 20. 2) The objective function of interest, such as minimizing the root mean squared error or minimizing the maximum variance of the queries.
Depending on the provided queries and objective function, the advertiser may aim to minimize the objective function within the constraints of a given privacy budget.
$\arg\min_{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}),\dots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{m};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}))$
[…]
Alternatively, the advertiser might have accuracy constraints for the queries and seek to minimize the total privacy budget. We use a constant vector $\mathbf{v}$ to specify variance constraints for queries.
[…]
Then we can solve for $\sigma_{1},\sigma_{2},\cdots,\sigma_{n}$ as the initial solution. The following theorem shows that using initial scales derived above, Mechanism 1 satisfies zCDP.
@@ 4.3 Initialize the Scales
Let $\mathbf{Q}=[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{2};\cdots;{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{m}]\in\mathbb{R}^{m\times n}$ be the input Query Workload matrix. For Mechanism 1, let $\rho_{1}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{budget}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$, $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the noise scales, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$, $\mathbf{L}=\mathbf{Q}\mathbf{B}^{-1}$ and $\mathbf{z}\sim N(0,\mathbf{I}_{n})$, then $\mathbf{L}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$ = $\mathbf{L}(\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{z})=\mathbf{Q}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+\mathbf{L}\mathbf{z}$ satisfies $\rho_{1}$-zCDP.
We next show an example of calculating the initial scales. Suppose on each day, the advertiser wants to get the total number of conversions up to today, the query ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}$ on day $i$ indicates a prefix sum function on ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}[t]=1$, when $1\leq t\leq i$, otherwise ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}[t]=0$.
The advertiser wants to minimize the weighted sum of variances given a fixed privacy budget. The loss function is the weighted sum of variances, $\gamma_{i}$ is the weight for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i}$.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}),\dots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{n};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}))=\sum_{i=1}^{n}\gamma_{i}^{2}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{i};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$
$=$ $\sum_{i=1}^{n}\left(\gamma_{i}^{2}\sum_{t=1}^{i}\sigma_{t}^{2}\right)=\sum_{i=1}^{n}\left(\sum_{j=i}^{n}\gamma_{j}^{2}\right)\sigma_{i}^{2}.$
The constraint on privacy budget is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{budget}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})\leq\rho$, or equivalently, $\Delta_{2}^{2}(f)\leq 2\rho$. Because we want to achieve $\rho$-zCDP while minimizing the total variance, the problem becomes,
$\min_{\sigma_{1}^{2},\cdots,\sigma_{n}^{2}}\quad$ $\left(\sum_{j=1}^{n}\gamma_{j}^{2}\right)\sigma_{1}^{2}+\left(\sum_{j=2}^{n}\gamma_{j}^{2}\right)\sigma_{2}^{2}+\cdots+\gamma_{n}^{2}\sigma_{n}^{2}$
$s.t.\quad$ $\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\frac{r_{2}^{2}}{\sigma_{2}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}\leq 2\rho.$ (7)
Problem (4.3) can be solved analytically using Cauchy inequality [51].
We give another example in the form of Problem 6. In this example, the advertiser has requirements for the accuracy of each query ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}$, the goal is to find a mechanism that has minimal privacy cost.
$\min_{\sigma_{1}^{2},\cdots,\sigma_{n}^{2}}\quad$ $\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\frac{r_{2}^{2}}{\sigma_{2}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}$ (8)
$s.t.\quad$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})\leq v_{j}\quad j=1,\cdots,m.$
As shown in Equation 2, the variance function is a linear combination of $\sigma_{i}^{2}$. Consequently, Problem 8 is a convex optimization problem that can be efficiently solved using any standard solver.
By selecting the initial scales, we maintain flexibility for different types of queries and objective functions. The key decision lies in determining these initial scales. We initialize the user contribution bound as $\bar{r}_{i}=1$ to compute the initial noise scale $\bar{\sigma}_{i}$. Once the actual user contribution bound $r_{i}$ is determined, we update the noise scales using $\sigma_{i}=\bar{\sigma}_{i}/\bar{r}_{i}*r_{i}$, as shown in Line 14 of Algorithm 1. This adjustment ensures that the privacy budget of the mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{budget}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$ remains unchanged. In the following sections, we describe how to adjust the noise scales based on the actual dataset to achieve more accurate measurements.
4.4 Estimate User Contribution Bound
Selecting an appropriate user contribution bound $r$ involves a trade-off between bias and variance. A small $r$ leads to the removal of numerous records, introducing substantial bias. Conversely, a large $r$ retains most records in the dataset but introduces significant noise into the results, resulting in high variance. To achieve a balance between bias and variance, we employ the differentially private quantile algorithm [22, 42] to estimate the contribution bound. Each day, we count the number of contributions per individual and select an appropriate private quantile as the user contribution bound for that day. The differentially private quantile is based on the exponential mechanism [35].
Theorem 4.4 (Exponential Mechanism [35, 38]).
Given utility function $u$ mapping (dataset, output) pairs to real valued scores with $L_{1}$ sensitivity $\Delta_{1}(u)=\max_{X\sim X^{\prime}}|u(X,o)-u(X^{\prime},o)|$, the Exponential Mechanism $M$ has output distribution $\mathcal{P}_{M}[M(X)=o]\propto\exp\left(\frac{\epsilon u(X,o)}{2\Delta_{1}(u)}\right)$. Here $\propto$ ignores the normalization factor. Then the mechanism $M$ is $\epsilon$-DP, and it satisfies $\frac{1}{8}\epsilon^{2}$-zCDP.
The utility function for the quantile is defined as
$u(X,o)=-||\{x\in X|x\leq o\}|-p*k|.$ (9)
For a given output $o$, |$|\{x\in X|x\leq o\}|$ is the count of records with values less than or equal to $o$. Here, $p$ denotes the quantile percentage, and $k$ is the total number of records. As demonstrated in [22, 42], the sensitivity of the utility function described above is 1. Algorithm 2 gives a detailed description of of PrivateQuantile. We can prove that the algorithm satisfies zCDP.
1:  Input: Dataset $D$; Quantile percentage $p$; privacy parameter $\epsilon$, bounding parameter $\Lambda$.
12:  Output: The selected user contribution bound based on $p$ quantile.
23:  $X$ = [].
4:  for Each user $u_{i}$ in $D$ do
5:     $c_{i}\leftarrow$ the number of records from $u_{i}$.
6:     $X$.append($c_{i}$).
37:  end for
8:  Sort $X$ in ascending order. $X=[c_{1},c_{2},\cdots,c_{k}]$.
49:  Replace $c_{i}>\Lambda$ with $\Lambda$. Define $c_{0}=0$ and $c_{k+1}$.
10:  for  $i=1$ to $k$  do
11:     Set $y_{i}=(c_{i+1}-c_{i})\exp(-\epsilon|i-p*k|/2)$.
12:  end for
13:  Sample an integer $i\in\{0,\cdots,k\}$ with probability $y_{i}/(\sum_{i=0}^{k}y_{i})$.
14:  $c\leftarrow$ uniformly sampled from $[c_{i},c_{i+1}]$.
15:  Return $c$.
Algorithm 2 PrivateQuantile